A water purification system for a rural business had an initial cost of $ 4000​, an expected life of 5​ years, and no scrap value. Use the​ double-declining-balance method of depreciation to find the book value at the end of 3 years.
Answer #1

Depreciation rate = (100% / 5) × 2 = 40%

Depreciation for first year = $4,000 × 40% = $1,600

Book value at the end of first year = $4,000 - $1,600 = $2,400

Depreciation for second year = $2,400 × 40% = $960

Book value at the end of second year = $2,400 - $960 = $1,440

Depreciation for third year = $1,440 × 40% = $576

Book value at the end of third year = $1,440 - $576 = $864